OpenStudy (anonymous):

The first two terms of a sequence are given as t1=2 and t2=4. The recusion formula is tn= (tn-1)^2-3tn-2. Determine the next four terms in the sequence

OpenStudy (anonymous):

do you understand what you have to do? \[t_n=t_{n-1}^2-3t_{n-2}\] so \[t_3=t_2^2-3t_1\]

OpenStudy (anonymous):

t3= (4)^2 - 3(2) =10, thank you:)

OpenStudy (anonymous):

yw